OpenStudy (anonymous):

A cheetah travels at a rate of 90 feet per second. • The distance d traveled by the cheetah is a function of seconds traveled t. Write a rule for the function. Hint: The distance d traveled by the cheetah is 90 times t (number of seconds). Put this wording into equation format. • How far will the cheetah travel in 25 seconds? Hint: Now substitute 25 in for t to get the distance traveled.

OpenStudy (perl):

a) d = 90 * t b) d = 90*25 = 2250